OpenStudy (anonymous):

Change the rectangular coordinates to polar coordinates with r > 0 and 0 ≤ θ ≤ 2π. a. (-1,1) b.(- sqroot 3,-1)

OpenStudy (anonymous):

a.) √2 b.) √-4

OpenStudy (anonymous):

whats the second part tho because it needs to be in (r,theta)

OpenStudy (anonymous):

I'm having trouble figuring out theta

OpenStudy (anonymous):

oh i know, hang on.

OpenStudy (anonymous):

tan-1(y/x)=tan-1(1/1)=tan-1(1)=45° so a.) (√2, 45°)

OpenStudy (anonymous):

yeah thats what i had but it says its wrong

OpenStudy (anonymous):

put it in radians 45°

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

pi/ 4

OpenStudy (anonymous):

i had (squroot 2, pi/4)

OpenStudy (anonymous):

oh wth

OpenStudy (anonymous):

hah, what said you were wrong?

OpenStudy (anonymous):

my homework is online

OpenStudy (anonymous):

try √2,3pi/4

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

do you happen to know the other one?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

√-4, 7pi/6 i believe

OpenStudy (anonymous):

was it right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how many tries do you have?

OpenStudy (anonymous):

100

OpenStudy (anonymous):

try √-4, pi/6

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i got it thanks for the help it was 2,7pi/6

OpenStudy (anonymous):

ha i got that just now, i just started learning this section in school.

OpenStudy (anonymous):

yeah same here..hate it

OpenStudy (anonymous):

ikr, i mean this is one of the easier sections.

OpenStudy (anonymous):

yeah i know i just hate math its soooo hard at Purdue

OpenStudy (anonymous):

i always did good in high school but college math is kicking my butt

OpenStudy (anonymous):

oh wow im a senior, i hate trig...its rediculous.

OpenStudy (anonymous):

yeah i took that my senior year also and I'm taking it again bc i only need so much math for Pre Vet and it just seems totally different from what i learned

OpenStudy (anonymous):

yeah, its hard to remember, and cool i wanted to be a vet, but theirs a lot of school, so im going to enroll in the navy after highschool.

OpenStudy (anonymous):

i cant really decide what i want to do now, thats why the navy.

OpenStudy (anonymous):

oh thats cool! good luck !! my boyfriend is in the marines

OpenStudy (anonymous):

nice. and thanks. (: